The critical angle for total internal reflection w
Answer #1


By using Snell's law

ng*sin(theta_g) = w*sin(theta_w)

ng*sin(theta_g) = nw*sin(90)


ng/nw = 1/sin(56) = 1.206


similarly

nw*sin(20) = ng*sin(theta)

sin(theta) = (nw/ng)*sin(20) = (1/1.206)*sin(20) = 0.283

theta = sin^(-1)(0.283) = 16.5 degrees